OpenStudy (cookiemonster18):

Can someone teach me how to solve this question ill fan and medal. 1. What is the ninth term of the arithmetic sequence defined by the rule below A(n)=-14(n-1)(2) A:232 B:230 C:2 D:4

OpenStudy (amistre64):

i think your 'rule' is not presented correctly

OpenStudy (cookiemonster18):

No it is. That's how it's shown on the worksheet A(n)=-14(n-1)(2)

OpenStudy (amistre64):

its just that that does not resemble any sort of arithmetic sequence that I am familiar with ...

OpenStudy (amistre64):

an arith seq is generated by some rule: A(n) = A(n-1) + d such that each new term is defined as the addition of d (the common difference) to the previous term.

OpenStudy (amistre64):

IF its a closed expression A(n) = A(1) + d(n-1) is equivalent

OpenStudy (amistre64):

if you are able to decipher your rule, then just let n=9 and work the math for the solution

OpenStudy (cookiemonster18):

C is correct it's 2 thanks @amistre64